OpenStudy (anonymous):

How do I tell if f(x)=2+1/x is invertible or not?

OpenStudy (loser66):

find it out

OpenStudy (anonymous):

Or any function for that matter, I just don't know the steps.

OpenStudy (loser66):

replace x = \(f^{-1}(x)\) and f(x) =x, what do you have?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

one way to do it is to go ahead and find the inverse

OpenStudy (misty1212):

if you can find it, it is invertible

OpenStudy (misty1212):

another way is to graph and see if it passes the horizontal line test

OpenStudy (anonymous):

replace x with x=2+1/y?

OpenStudy (misty1212):

another way it to write \[2+\frac{1}{a}=2+\frac{1}{b}\]and see if, by algebra, you can prove \(a=b\) that is the definition of a one to one function

OpenStudy (misty1212):

yeah that is how you find the inverse solve \[x=2+\frac{1}{y}\] for \(y\) should take only two steps

OpenStudy (anonymous):

Got y=1/(x-2)

OpenStudy (misty1212):

me too